OpenStudy (anonymous):

Find the least common multiple for the given pair. x² - 25 and x² + 10x + 25

OpenStudy (anonymous):

@ganibl can you help me?

OpenStudy (anonymous):

x^2 + 10x + 25 x^2 + 5x + 5x + 25 x(x + 5) + 5 (x + 5) (x + 5)^2

OpenStudy (anonymous):

x^2 - 35 => (x + 5) ( x - 5) Since (a+b)(a - b) = a^2 - b^2

OpenStudy (anonymous):

12x^2y^3 and 14x^3y^2

OpenStudy (anonymous):

Sorry its x^2 - 25 I wrote as x^2 - 35

OpenStudy (anonymous):

Thats easy just eliminate x^2 y^2 from numerator and denominator and see what u get

OpenStudy (anonymous):

what do you mean from numerator

OpenStudy (anonymous):

Is it 12x^2y^3/14x^3y^2???

OpenStudy (anonymous):

no, it is Find the least common multiple for the given pair. 12x^2y^3 and 14x^3y^2

OpenStudy (anonymous):

oh ok It should be 2x^2y^2 then

OpenStudy (anonymous):

Thats the least common factor which multiplies the given terms :)

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

Last one! Find the least common multiple for the given pair. 16x^2 – 25 and 4x^2 – x - 5

OpenStudy (anonymous):

@ganibl

OpenStudy (anonymous):

16x^2 - 25 can be written as (4x + 5) (4x - 5) since (a+b)(a - b) = a^2 - b^2 4x^2 -x - 5 we need to simplify this 4x^2 -5x + 4x -5 x(4x - 5) + 1 (4x - 5) (4x - 5) (x + 1)

OpenStudy (anonymous):

So from both the common factor is 4x - 5

OpenStudy (anonymous):

So 4x^2-5?

OpenStudy (anonymous):

or just 4x-5?

OpenStudy (anonymous):

Its 4x - 5

OpenStudy (anonymous):

Okay! thank you for your help!